# Engagement Maximization

### v5

@@ A.1 Proof of Proposition 1
$\mathbb{E}^{\pi^{\prime}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi^{*}(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)\right]-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{1})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{1})\right]\right]$
$+\mathbb{E}^{\pi^{*}(q_{1})}\left[\hat{u}(q_{1})-\frac{\bar{\kappa}}{\chi}H(q_{1})-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$\geq$ $0$
Therefore, $\pi^{\prime}$ is feasible and strictly improves $\pi^{*}$, contradiction. ∎
A.2 Proof of Proposition 2
@@
Let $P$ be the law of the process defined in (5), and define the stopping time $\tau=\inf\left\{t\in\mathbb{R}_{+}|q_{t}\neq q_{0}\right\}$. Evidently, conditional on continuation, $\lim_{h\downarrow 0}\frac{1}{h}\mathbb{E}^{P}[H(q_{t})-H(q_{(t-h)^{-}})|\mathcal{F}_{(t-h)^{-}}]=\lim_{h\downarrow 0}\frac{1}{h}(1-e^{-\frac{\chi\cdot h}{\mathbb{E}^{\pi}[H(q)-H(q_{0})]}})(\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})])=\chi$. Thus, the information constraint is satisfied. Since $q_{\tau}\sim\pi$, the principal’s utility is $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$. What remains to be verified is the agent’s optimality condition.
@@
$\mathbb{E}^{P}[\widehat{\mu}(q_{\tau^{\prime}})-\bar{\kappa}\tau^{\prime}|\mathcal{F}_{0}]=$ $\mathrm{Prob}(\tau^{\prime}<\tau)\left(\widehat{\mu}(q_{0})-\bar{\kappa}\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau)\left(\mathbb{E}^{P}\left[\widehat{\mu}(q_{\tau})-\bar{\kappa}\tau^{\prime}|\tau^{\prime}\geq\tau\right]\right)$
$\leq$ $\mathrm{Prob}(\tau^{\prime}<\tau)\left(\mathbb{E}^{\pi}[\widehat{\mu}(q)-\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]-\bar{\kappa}\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau)\left(\mathbb{E}^{\pi}[\widehat{\mu}(q)]-\bar{\kappa}\mathbb{E}^{P}\left[\tau|\tau^{\prime}\geq\tau\right]\right)$
$=$ $\mathrm{Prob}(\tau^{\prime}<\tau)\left(\mathbb{E}^{\pi}[\widehat{\mu}(q)-\bar{\kappa}\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau)\left(\mathbb{E}^{\pi}[\widehat{\mu}(q)]-\bar{\kappa}\mathbb{E}^{P}\left[\tau|\tau^{\prime}\geq\tau\right]\right)$
$=$ $\mathbb{E}^{P}[\widehat{\mu}(q)-\bar{\kappa}\tau|\mathcal{F}_{0}].$
The first equality is from the definition of conditional expectations and the process defining $q_{t}$, (5). The first inequality is from the constraint $\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$ in the relaxed problem (R) and $E[\tau|\tau^{\prime}\geq\tau]\leq E[\tau^{\prime}|\tau^{\prime}\geq\tau]$. The second equality is from the memorylessness property of $\tau$, [^24] Specifically, $\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau-\tau^{\prime}|\tau^{\prime}<\tau\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau|\mathcal{F}_{0}\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]$. and the last from the definition of conditional expectations. Note that if the constraint in (R) is slack, the first inequality is strict if $\tau^{\prime}$ is not equal to $\tau$ $P$-a.e. In this case, the $\alpha$-dilution of $\pi$ strongly implements principal’s utility level $\rho\mathbb{E}^{\pi}[G(q_{t})-G(\bar{q}_{0})]$. ∎
A.3 Proof of 2
@@
If $V^{B}(q)>\hat{u}(q)$, then let $\pi$ implement $V^{B}(q)$. Let $\pi^{\prime}$ satisfies $\mathbb{E}^{\pi}[H(q^{\prime})-H(q)]=\mathbb{E}^{\pi^{\prime}}[H(q^{\prime})-H(q)]$ and $\mathbb{E}^{\pi^{\prime}}[G(q^{\prime})-G(q)]>0$. Then, the dilution of a convex combination of $\pi$ and $\pi^{\prime}$ is feasible, incentive compatible for sufficiently small weight on $\pi^{\prime}$ and strictly profitable for the principal. Therefore, $J(q)>0$.
Next, suppose for the purpose of contradiction that $\pi$ implies $V^{B}(\bar{q}_{0})$ but $\mathbb{E}^{\pi}[V^{B}(q)-\hat{\mu}(q)]>0$. For each $q\in\text{Supp}(\pi)$, let $\pi_{q}$ be the probability measure that implements $V^{B}(q)$. Let $\pi^{\prime}=\mathbb{E}^{\pi(q)}[\pi^{q}]$.
$\mathbb{E}^{\pi^{\prime}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)\right]-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{1})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{1})\right]\right]$
$+\mathbb{E}^{\pi(q_{1})}\left[\hat{u}(q_{1})-\frac{\bar{\kappa}}{\chi}H(q_{1})-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$>$ $V^{B}(\bar{q}_{0})-\hat{u}(\bar{q}_{0}).$
This contradicts the optimality of $\pi$. ∎
A.4 Proof of Proposition 3
@@
Observe first by Proposition 1 that $\bar{J}(\bar{q}_{0})>0$ implies $\lambda>0$ (where $\lambda$ is defined as in that proposition).
@@
$U^{p}(q)=\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)+\lambda G(q).$
Define $H^{*}(q)$ as the concave envelope of $\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$. Observe that any
$\pi^{*}\in\arg\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)+\lambda G(q)\right]$
places full support on points where $H^{*}(q)=\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$, by the convexity of $\hat{u}(q)$, and for each $q^{\prime}\in\mathcal{P}(X)$ there exists a $\pi^{\prime}\in\mathcal{P}(\mathcal{P}(X))$ with $\mathbb{E}^{\pi^{\prime}}[q]=q^{\prime}$ and
$\hat{u}(q^{\prime})-H^{*}(q^{\prime})\leq\mathbb{E}^{\pi^{\prime}}[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)+\lambda G(q)],$
@@
$\pi^{*}\in\arg\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}\left[\hat{u}(q)-H^{*}(q)\right].$
Now observe that $q^{*}\in Q^{p}(\bar{q}_{0})$ and $\bar{J}(\bar{q}_{0})>0$ implies by Proposition 1 and Theorem 1 that there exists a $\pi^{*}\in\mathcal{P}(\mathcal{P}(X))$ with $q^{*}\in\mathrm{Supp}\pi^{*}$ such that $\pi^{*}$ is a solution to a static rational inattention problem. By the Lagrangian lemma of Caplin et al. [2022] applied to the problem with the cost function $H^{*}$, there exists a linear function $L^{p}:\mathcal{P}(X)\rightarrow\mathbb{R}$ such that
$\hat{u}(q)-H^{*}(q)-L^{p}(q)\leq\hat{u}(q^{*})-H^{*}(q^{*})-L^{p}(q^{*})$
with equality for all $q\in\mathrm{Supp}\pi^{*}$. By $H^{*}(q)\geq\frac{\bar{\kappa}}{\chi}H(q)-\lambda G(q)$, with equality for all $q\in\mathrm{Supp}\pi^{*}$,
@@
Note that $\hat{u}$ is piecewise linear and convex and both $H$ and $G$ are continuously differentiable. It follows that $\hat{u}$ can not involve a convex kink at any $q\in\mathrm{Supp}\pi^{*}$, and hence if such a $q$ lies in the relative interior of $\mathcal{P}(X)$,
@@
but this requires (by $\bar{J}(\bar{q}_{0})=E^{\pi^{*}}[G(q)]-G(\bar{q}_{0})]>0$) that
@@ A.5 Proof of 4
Per Proposition 1, the optimal $\pi$ necessarily leads to $J(\bar{q}_{0})=\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$. Then, (R) reduces to:
$\sup_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=q_{0}\end{subarray}}$ $\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$
$s.t.$ $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]\leq\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})].$
$\forall\pi\in\mathcal{P}(\mathcal{P}(X))$, let $\pi_{1}=\pi|_{X_{1}}$, $\pi_{2}=\delta_{\bar{q}_{0}|_{X_{2}}}$ and $\pi^{\prime}=\pi_{1}\otimes\pi_{2}$. Note that $\pi$ is Blackwell more informative than $\pi^{\prime}$. Then,
$\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]=\mathbb{E}^{\pi_{1}}[\widehat{\mu}(q,\bar{q}_{0}|_{X_{2}})-\widehat{\mu}(\bar{q}_{0})]=\mathbb{E}^{\pi^{\prime}}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})];$
$\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]\geq\mathbb{E}^{\pi^{\prime}}[H(q)-H(\bar{q}_{0})].$
Therefore, $\pi^{\prime}$ is feasible and weakly improves $\pi$. ∎
A.6 Proof of 5

### v6

@@ A.2 Proof of Proposition 1
$\mathbb{E}^{\pi^{\prime}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi^{*}(q)}\left[\mathbb{E}^{\pi^{q}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})\right]-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi^{*}(q)}\left[\mathbb{E}^{\pi^{q}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})-\widehat{u}(q)+\frac{\kappa}{\chi}H(q)\right]\right]$
$+\mathbb{E}^{\pi^{*}(q)}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]$
$\geq$ $0.$
Therefore, $\pi^{\prime}$ is feasible in (R) and strictly improves on $\pi^{*}$, contradicting the optimality of $\pi^{*}$. It follows by the non-negativity of $J(q)$ that $\mathbb{E}^{\pi^{*}}[J(q)]=0$. ∎
A.3 Proof of Proposition 2
@@
Under the $\alpha$-dilution policy, if $t<\tau$,
$\mathbb{E}^{P}[H(q_{t+\delta})-\chi(t+\delta)|\mathcal{F}_{t},\tau>t]$ $=(1-e^{-\alpha\delta})\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]+H(\bar{q}_{0})-\chi(t+\delta).$
Using $\alpha=\chi(\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})])^{-1}$ and $x\geq 1-e^{-x}$,
$\mathbb{E}^{P}[H(q_{t+\delta})-\chi(t+\delta)|\mathcal{F}_{t},\tau>t]\leq H(\bar{q}_{0})-\chi(t).$
as required by (1). That constraint holds by assumption if $t\geq\tau$.
By $q_{\tau}\sim\pi$, the principal’s utility is $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$. What remains to be verified is the agent’s optimality condition.
@@
$\mathbb{E}^{P}[\widehat{u}(q_{\tau^{\prime}})-\kappa\tau^{\prime}|\mathcal{F}_{0}]=$ $\mathrm{Prob}(\tau^{\prime}<\tau|\mathcal{F}_{0})\left(\widehat{u}(\bar{q}_{0})-\kappa\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau|\mathcal{F}_{0})\left(\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-\kappa\tau^{\prime}|\tau^{\prime}\geq\tau,\mathcal{F}_{0}\right]\right)$
$\leq$ $\mathrm{Prob}(\tau^{\prime}<\tau|\mathcal{F}_{0})\left(\mathbb{E}^{\pi}[\widehat{u}(q)]-\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]-\kappa\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau|\mathcal{F}_{0})\left(\mathbb{E}^{\pi}[\widehat{u}(q)]-\kappa\mathbb{E}^{P}\left[\tau|\tau^{\prime}\geq\tau,\mathcal{F}_{0}\right]\right)$
$=$ $\mathrm{Prob}(\tau^{\prime}<\tau|\mathcal{F}_{0})\left(\mathbb{E}^{\pi}[\widehat{u}(q)-\kappa\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]\right)$
$+\mathrm{Prob}(\tau^{\prime}\geq\tau|\mathcal{F}_{0})\left(\mathbb{E}^{\pi}[\widehat{u}(q)]-\kappa\mathbb{E}^{P}\left[\tau|\tau^{\prime}\geq\tau,\mathcal{F}_{0}\right]\right)$
$=$ $\mathbb{E}^{P}[\widehat{u}(q)-\kappa\tau|\mathcal{F}_{0}].$
The first equality is from the definition of conditional expectations, the process defining $q_{t}$, equation (2), and the convention that $q_{t}=q_{\tau}$ for $t>\tau$. The first inequality is from the constraint $\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]\geq\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$ in the relaxed problem (R) and $E[\tau|\tau^{\prime}\geq\tau]\leq E[\tau^{\prime}|\tau^{\prime}\geq\tau]$. The second equality is from the memorylessness property of $\tau$, [^24] Specifically, $\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]=\mathbb{E}^{P}\left[\tau-\tau^{\prime}|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]=\mathbb{E}^{P}\left[\tau|\mathcal{F}_{0}\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau,\mathcal{F}_{0}\right]$. and the last from the definition of conditional expectations. Note that if the constraint in (R) is slack, the first inequality is strict if $\tau^{\prime}$ is not equal to $\tau$ $P$-a.e. In this case, the $\alpha$-dilution of $\pi$ strongly implements principal’s utility level $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$. ∎
A.4 Proof of ˜2
@@
If $V^{B}(\bar{q}_{0})>\widehat{u}(\bar{q}_{0})$, then let $\pi$ implement $V^{B}(\bar{q}_{0})$. By the assumption that full information is strictly beneficial, let $\pi^{\prime}$ be the distribution of posterior beliefs corresponding to full information and note by Assumption $\mathbb{E}^{\pi^{\prime}}[G(q^{\prime})-G(\bar{q}_{0})]>0$. Then, the dilution of a convex combination of $\pi$ and $\pi^{\prime}$ is feasible, incentive compatible for sufficiently small weight on $\pi^{\prime}$ and strictly profitable for the principal. Therefore, $J(\bar{q}_{0})>0$.
Next, suppose for the purpose of contradiction that $\mathbb{E}^{\pi_{A}}[V^{B}(q)-\widehat{u}(q)]>0$. For each $q\in\text{Supp}(\pi_{A})$, let $\pi^{q}$ be the probability measure that implements $V^{B}(q)$. Define $\pi^{\prime}\in\Pi(\bar{q}_{0})$ by $d\pi^{\prime}(q^{\prime})=\mathbb{E}^{\pi_{A}(q)}[d\pi^{q}(q^{\prime})]$. Then
$\mathbb{E}^{\pi^{\prime}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi_{A}(q)}\left[\mathbb{E}^{\pi^{q}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})\right]-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi_{A}(q)}\left[\mathbb{E}^{\pi^{q}(q^{\prime})}\left[\widehat{u}(q^{\prime})-\frac{\kappa}{\chi}H(q^{\prime})-\widehat{u}(q)+\frac{\kappa}{\chi}H(q)\right]\right]$
$+\mathbb{E}^{\pi_{A}(q)}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)-\widehat{u}(\bar{q}_{0})+\frac{\kappa}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi_{A}(q)}\left[V^{B}(q)-\hat{u}(q)\right]+V^{B}(\bar{q}_{0})-\widehat{u}(\bar{q}_{0})$
$>$ $V^{B}(\bar{q}_{0})-\widehat{u}(\bar{q}_{0}).$
This contradicts the optimality of $\pi_{A}$, and by $V^{B}(q)\geq\hat{u}(q)$ we must have $\mathbb{E}^{\pi_{A}}[V^{B}(q)-\hat{u}(q)]=0$. ∎
A.5 Proof of Proposition 3
@@
Observe first by Lemma 2 that $V^{B}(\bar{q}_{0})>\widehat{u}(\bar{q}_{0})$ implies $\lambda>0$ (where $\lambda$ is defined as in Proposition 1).
@@
$U^{p}(q)=\widehat{u}(q)-\frac{\kappa}{\chi}H(q)+\lambda G(q).$
Define $H^{*}(q)$ as the concave envelope of $\frac{\kappa}{\chi}H(q)-\lambda G(q)$. Observe that any
$\pi^{*}\in\arg\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)+\lambda G(q)\right]$
places full support on points where $H^{*}(q)=\frac{\kappa}{\chi}H(q)-\lambda G(q)$, by the convexity of $\widehat{u}(q)$, and for each $q^{\prime}\in\mathcal{P}(X)$ there exists a $\pi^{\prime}\in\mathcal{P}(\mathcal{P}(X))$ with $\mathbb{E}^{\pi^{\prime}}[q]=q^{\prime}$ and
$\widehat{u}(q^{\prime})-H^{*}(q^{\prime})\leq\mathbb{E}^{\pi^{\prime}}[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)+\lambda G(q)],$
@@
$\pi^{*}\in\arg\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}\left[\widehat{u}(q)-H^{*}(q)\right].$
Now observe that $q^{*}\in Q^{p}(\bar{q}_{0})$ and $J(\bar{q}_{0})>0$ implies by Proposition 1 and Theorem 1 that there exists a $\pi^{*}\in\mathcal{P}(\mathcal{P}(X))$ with $q^{*}\in\mathrm{Supp}\pi^{*}$ such that $\pi^{*}$ is a solution to a static rational inattention problem. By the Lagrangian lemma of caplin2017shannon applied to the problem with the cost function $H^{*}$, there exists a linear function $L^{p}:\mathcal{P}(X)\rightarrow\mathbb{R}$ such that
$\widehat{u}(q)-H^{*}(q)-L^{p}(q)\leq\widehat{u}(q^{*})-H^{*}(q^{*})-L^{p}(q^{*})$
with equality for all $q\in\mathrm{Supp}\pi^{*}$. By $H^{*}(q)\geq\frac{\kappa}{\chi}H(q)-\lambda G(q)$, with equality for all $q\in\mathrm{Supp}\pi^{*}$,
@@
Note that $\widehat{u}$ is piecewise linear and convex and both $H$ and $G$ are continuously differentiable. It follows that $\widehat{u}$ can not involve a convex kink at any $q\in\mathrm{Supp}\pi^{*}$, and hence if such a $q$ lies in the relative interior of $\mathcal{P}(X)$,
@@
but this requires (by $J(\bar{q}_{0})=E^{\pi^{*}}[G(q)]-G(\bar{q}_{0})]>0$) that
@@ A.6 Proof of ˜4
Per Proposition 1, the optimal $\pi$ necessarily leads to $J(\bar{q}_{0})=\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]$. Then, (R) reduces to:
$\sup_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=q_{0}\end{subarray}}$ $\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]$
$s.t.\$$\frac{\kappa}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]\leq\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})].$
Proposition 1 implies that $\pi$ could be picked with finite support. Let $s$ be the signal that induces $\pi$ per the Bayes rule. Let $p(s|x_{1},x_{2})$ denote the conditional distribution of $s$. Let $s^{\prime}$ be another signal that has conditional distribution $p(s^{\prime}=z|x_{1},x_{2})=\mathbb{E}^{\bar{q}_{0}}[p(s=z|x_{1},x_{2})|x_{1}]$, i.e., $s^{\prime}$ contains $s$’s information about $x_{1}$ only. Let $\pi^{\prime}$ be the distribution of posterior induced by $s^{\prime}$. Evidently, since $u$ depends on $x_{1}$ only,
$\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi_{1}}[\widehat{u}(q,\bar{q}_{0}|_{X_{2}})-\widehat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi^{\prime}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})].$
Therefore, $\pi^{\prime}$ is feasible. Next, we show that $\pi^{\prime}$ weakly improves $\pi$. Note that
$\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]=I(s;x_{1},x_{2})=\mathbb{E}^{\bar{q}_{0}}[D_{KL}(p(s|x_{1},x_{2}))||p(s)];$
$\mathbb{E}^{\pi^{\prime}}[H(q)-H(\bar{q}_{0})]=I(s^{\prime};x_{1},x_{2})=\mathbb{E}^{\bar{q}_{0}}[D_{KL}(p(s^{\prime}|x_{1}))||p(s^{\prime})]$
$\implies$ $\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]-\mathbb{E}^{\pi^{\prime}}[H(q)-H(\bar{q}_{0})]$
$=\mathbb{E}^{\bar{q}_{0}}\left[\log\frac{p(s|x_{1},x_{2})}{p(s)}-\log\frac{p(s|x_{1})}{p(s)}\right]$
$=\mathbb{E}^{\bar{q}_{0}}[D_{KL}(p(s|x_{1},x_{2})||p(s|x_{1}))]\geq 0,$
strictly if $p(s|x_{1},x_{2})\neq p(s|x_{1})$ for some $(x_{1},x_{2})$ in the support of $\bar{q}_{0}$.
A.7 Proof of ˜5